# Search and Rescue on a poset

### 08 Dec 2023 v1

@@ 1. Introduction
SR games were first defined by Lidbetter in [14] and quickly proved to be a fruitful avenue of new study. It has been applied to scheduling problems [1, 2], object detection [4], rendezvous problems [13], and adaptive search [15]. For such applications it is natural to impose an order on the set of locations, as some jobs can be performed only if other jobs have finished. In our paper, we therefore consider SR games on a partially ordered set $X$. A search now has to respect the order and is only allowed if $\sigma(i)<\sigma(j)$ implies $i<j$. For instance, if $m\in X$ is a maximal element for which $x<m$ for all $x\not=m$, then a search halts in $m$. We call this an ordered search-and-rescue game, or simply an OSR game. We also consider the stronger restriction that $\sigma$ is only allowed if $i<j$ implies $\sigma(i)<\sigma(j)$. A function with this property is called a chain  [21] and we say that this game is a chained SR game, or simply a CSR game. If we represent the partial order by a Hasse diagram, then we get a search game on a (directed) graph, which is a topic of its own [10]. A search is a path on the Hasse diagram in which Searcher may skip some vertices but is not allowed to backtrack. She can either search a node or skip it without turning back. This appears to be a new game on a graph.
@@ 4. Correlated Search and Rescue
If the weights are not $\leq 1$ then depth-first search need not be optimal and we were unable to find a general solution for this game. We conjecture that the optimal search is backjumping [25] in the opposite case when all factors are $\geq 1$. Under backjumping, once the search reaches a leaf of the tree, it continues at an unsearched node that is closest to the root. In this respect it is the opposite of depth-first-search which continues at the unsearched node that is closest to the leaf.
Conjecture 13.
Optimal search in the OSR game for a completely reducible distribution on an unordered $X$ is backjumping if all Bayesian factors are $\geq 1.$
To specify this conjecture, suppose that a node $x$ has two children $a$ and $b$ with offspring $A$ and $B$. If a search first visits a leaf in $A$ then it does not re-enter $A$ before visiting a leaf of $B$. The backjump is maximal. Backjump search is well-studied in AI and automatic theorem proving [19].
@@
This bound is equal to the upper bound for uncorrelated random variables that we found in lemma 2. It is possible to extend that lemma to correlated Bernoulli random variables, under a restriction on their distribution.
@@ 6. Conclusion
We have presented a new type of Search and Rescue game on partially ordered sets. We were able to solve the game in polynomial time in terms of the number of locations for uncorrelated random variables. We showed that the SR game on a tree can be interpreted as a game with correlated random variables, for which the distribution is completely reducible. This game was solved by Lidbetter under the restriction that Bayesian factors are bounded by one. The solution of the game for general Bayesian factors remains open, but we did find general bounds on the value of the game, and we conjecture that optimal search is backjumping if all factors are at least equal to one. We hope that our extension of SR games to posets and Bayesian networks provides a fruitful avenue of further research and applications.

### 13 Mar 2024 v2

@@ 1. Introduction
SR games were first defined by Lidbetter in [14] and quickly proved to be a fruitful avenue of new study. They have been applied to scheduling problems [1, 2], object detection [4], rendezvous problems [13], and adaptive search [15]. For such applications it is natural to impose an order on the set of locations, as some jobs can be performed only if other jobs have finished. In our paper, we therefore consider SR games on a partially ordered set $X$. A search now has to respect the order and is only allowed if $\sigma(i)<\sigma(j)$ implies $i<j$. For instance, if $m\in X$ is a maximal element for which $x<m$ for all $x\not=m$, then a search halts in $m$. We call this an ordered search-and-rescue game, or simply an OSR game. We also consider the stronger restriction that $\sigma$ is only allowed if $i<j$ implies $\sigma(i)<\sigma(j)$. A function with this property is called a chain  [21] and we say that this game is a chained SR game, or simply a CSR game. If we represent the partial order by a Hasse diagram, then we get a search game on a (directed) graph, which is a topic of its own [10]. A search is a path on the Hasse diagram in which Searcher may skip some vertices but is not allowed to backtrack. She can either search a node or skip it without turning back. This appears to be a new game on a graph.
@@ 4. Correlated Search and Rescue
Lidbetter showed that depth-first search is optimal. Searcher starts at the root and continues until it reaches a leaf and backtracks, continuing at an unsearched node that is closes to the leaf. What can be said if all Bayesian factors are $\geq 1$? A search is backjumping [25] if it continues at an unsearched node that is closest to the root. It is a common method in AI and automatic theorem proving [19] to prevent a search more dispersive. To specify this, suppose that a node $x$ has two children $a$ and $b$ with offspring $A$ and $B$. If a search first visits a leaf in $A$ then it does not re-enter $A$ before visiting a leaf of $B$. Backjumping search is the opposite of backtracking search It is a natural procedure if all Bayesian factors are $\geq 1$, when it is unlikely that a search of a nearby leaf will be successful. Numerical experiments show that it is not true that the optimal search is a mix of backjumping searches if all Bayesian factors are $\geq 1$. They do indicate that the following weaker statement is true.
Conjecture 14.
Consider the OSR game with a completely reducible distribution with all Bayesian factors $\geq 1$. An optimal Searcher strategy contains a pure strategy that is backjumping.
@@ 6. Conclusion
We have presented a new type of Search and Rescue game on partially ordered sets. We were able to solve the game in polynomial time in terms of the number of locations for uncorrelated random variables. We showed that the SR game on a tree can be interpreted as a game with correlated random variables, for which the distribution is completely reducible. This game was solved by Lidbetter under the restriction that Bayesian factors are bounded by one from above. If they are bounded by one from below, the solution appears to be more difficult. We conjecture that an optimal search involves a backjumping search. A full solution of the game for general Bayesian factors seems to be out of reach, but we did find general bounds on the value of the game. We hope that SR games on posets and Bayesian networks provide a fruitful avenue of further research and applications.